S8= 2250, r= 2, A3= ?
Answer by stanbon(75887)

S8= 2250, r= 2, A3= ?
---------------------
Assuming your S(8) is the sum of the 1st 8 elements of the series,
S(8) = a[(1-2^8)/(1-2)] = 2250
a[(-255/-1)] = 2250
a = 2250/255 = 8.82
-----
a(3) = a*r^2 = 8.82*4 = 35.29
-------------
Cheers,
Stan H.
-------------